How long will it take Brian to increase his velocity from 5.0 m/s to 9.0 m/s if his rate of acceleration is 0.35 m/s 2
1
Expert's answer
2020-06-02T17:51:20-0400

We can find the time needed to increase Brian's velocity from 5.0 m/s to 9.0 m/s from the kinematic equation:


"v = v_0 + at,""t = \dfrac{v - v_0}{a} = \dfrac{9.0 \ \dfrac{m}{s} - 5.0 \ \dfrac{m}{s}}{0.35 \ \dfrac{m}{s^2}} = 11.43 \ s."

Answer:

"t = 11.43 \ s."